OpenStudy (anonymous):

f(x) = 4x^2 + 5 Find the value of f(x) when x = 2

OpenStudy (dmndlife24):

f(x) = 4x^2 + 5 f(2) = 4(2)^2 + 5 = 16 + 5 = ? (add 16 and 5)

OpenStudy (anonymous):

The answer is 21

OpenStudy (anonymous):

Thank you both so much. :)